# Adaptive Orchestration for Inference of Large Foundation Models at the Edge

### v1

@@ Abstract
Large Foundation Models (LFMs), including multi-modal and generative AI models, promise to unlock new capabilities for next-generation Edge AI applications. However, performing inference with LFMs in resource-constrained and heterogeneous edge environments presents significant challenges for workload orchestration. We propose a novel adaptive orchestration method and system tailored specifically for managing distributed inference workloads across multi-access edge computing (MEC) infrastructures. Our approach enhances traditional workload orchestration by introducing dynamic methods including: (1) adaptive workload distribution that selects optimal, inter-connected edge nodes based on runtime capacity profiling, (2) dynamic redistribution of LFM partitions as operational conditions evolve, and (3) real-time reconfiguration (e.g., re-splitting) of LFM layers to balance performance and privacy requirements. Specifically, our proposed framework introduces an architecture for adaptive split inference, enabling real-time, QoS-aware management of inference workloads. We present a reference architecture, detail operational mechanisms, and demonstrate its application through various use cases in real-world scenarios.
Index Terms:
Foundation Models, Split Inference, Multi-Access Edge Computing, 6G, Edge AI
I Introduction
@@
Distributed Split Inference [1, 2] strategies have typically employed static splits of inference workloads, where some computation is executed locally, while heavier computation tasks are outsourced, alleviating the computational burden at the client device. Such splits are mostly predetermined before execution and thus lack adaptability to dynamic and heterogeneous operational conditions, such as fluctuating network reliability, changing node utilization, or intermittent connectivity. Consequently, these approaches produce suboptimal performance, compromising latency, resource utilization, and quality of service (QoS) guarantees, especially in mission-critical or latency-sensitive applications such as those found in financial services, industrial manufacturing, retail operations, and logistics. This problem becomes even more acute in resource-constrained and heterogeneous edge environments, where multiple users rely on accessing shared edge compute resources, and where data privacy regulations (e.g., GDPR [3]) often restrict offloading computations to the cloud.
@@
We introduce an adaptive split inference orchestration framework, extending existing workload orchestration systems with domain-specific capabilities that are specifically tailored for LFMs, such as (multi-modal) LLMs. We introduce the following capabilities by leveraging the modular architecture of these models:
@@
    Redistribution of split model partitions across connected edge nodes to dynamically optimize resources under changing conditions.
@@
As our solution emphasizes split inference. However privacy can be implemented as an additional feature at no cost, when sensitive LFM layers can be executed locally, which makes reverse engineering data from model weights significantly more challenging for attackers. Through this approach, we establish a foundation for privacy-preserving, real-time, QoS-aware AI inference in edge networks, aligning with key 6G objectives of seamless connectivity, low inference latency, and intelligent edge resource management.
This work contributes to the state-of-the-art by providing:
  • A reference architecture for adaptive distributed inference orchestration in Edge AI.
  • A method for dynamic workload redistribution that allows foundation models to operate efficiently under varying computational workloads, intermittent connectivity, and network variability.
  • An approach to enhance the capability for real-time QoS-aware model partitioning, dynamically optimizing inference task execution to ensure service level agreements (SLAs) and Quality-of-Service targets.
  • An approach to enhanced privacy preservation in LFM inference, ensuring sensitive data and computations remain localized to Edge nodes, thus reducing compliance risks.
II Background
II-A Large Foundation Models at the Edge
@@
Next-generation 6G-enabled services in dense urban environments will need to support a multitude of AI-driven applications underpinned by Large Foundation Models (LFMs) [4, 5]. However, deploying such large models typically requires significant computational resources and raises privacy concerns when handling sensitive data, making their adoption particularly challenging for inference in edge environments [6, 7]. Further, as organizations strive to keep data on-premise (e.g., for regulatory compliance such as GDPR [3]), optimizing distributed inference becomes a critical enabler of low-latency and privacy-preserving AI services [8, 9].
@@
For that, next-generation networks make use of Multi-Access Edge Computing (MEC) infrastructures [10], which embed computing resources directly within the network. While this brings compute closer to end-users, a single MEC-enabled base station can quickly become saturated by various inference workloads such as from smart city and crowd management, personalized user applications, or industrial applications [11]. Looking ahead, 6G is envisaged to evolve beyond a mere network infrastructure upgrade into intelligent workload orchestration systems, enabling distributed, LFM-based AI services that seamlessly shift computation across user devices, edge, and cloud as network and compute conditions change [12, 13].
II-B Challenges in AI Workload Orchestration
@@
The current industry norm has been to integrate general-purpose orchestration platforms (e.g., Kubernetes or native MEC orchestrators), which facilitate application deployment and scaling but were not designed for challenges in inference scaling of modern LFM architectures, including fine-grained model partitioning and real-time, potentially hardware-accelerated inference optimization [14]. Existing solutions thus primarily target stateless services or relatively simple microservices, neglecting the unique requirements of edge-based AI pipelines, particularly for dynamic model splitting, QoS-driven scheduling, and adaptive resource reallocation [15].
@@
Nonetheless, in Edge AI scenarios involving LFMs, traditional workload orchestration platforms fall short of meeting edge AI-centric inference demands as they lack mechanisms to dynamically redistribute or reconfigure large model partitions based on real-time changes in network conditions, node utilization, or connectivity [16]. As a result:
@@
  • Resource utilization becomes imbalanced, either overloading certain nodes or leaving others underutilized, leading to missed Service-Level Agreements (SLAs).
@@
II-C Distributed and Adaptive Split Inference
To alleviate computational demands, distributed split inference (DSI) [20] has emerged as an approach within Edge AI, which partitions a model across different compute locations (e.g., client device, MEC edge, cloud) to balance on-device processing with remote offloading. Here, early, lightweight, or privacy-sensitive model layers are often chosen to be executed locally on-device or on trusted MEC nodes, extracting compact feature maps to reduce data transmission overhead. Subsequent layers are then offloaded to remote cloud environments and processed by more powerful servers [20, 21].
@@ II-C Distributed and Adaptive Split Inference
Thus, in standard implementations, orchestrators cannot dynamically decide to offload additional LFM layers to another edge node or re-split the network, leaving significant performance and reliability gains unrealized. Note that this problem does not originate from within the network itself as current 5G and future 6G architectures natively implement adaptive resource management strategies for various services and network slices [24]. In current deployments, effectiveness thus remains limited by the absence of real-time dynamic orchestration policies tailored explicitly to modern AI workloads and complex LFM deployment scenarios.
This is especially problematic for heterogeneous compute nodes, which complicates uniform deployment strategies. Thus, a one-size-fits-all static partitioning rarely works, as local workloads, performance constraints, and available resources differ from one site to another [6].
II-D Key Design Goals for Adaptive LFM Split Inference
@@ III Proposal
III-A Reference Architecture
Figure 1: Reference architecture of the proposed adaptive split inference orchestration. Sub-split models (S1, S2, S3) are deployed across edge/cloud nodes, while a central orchestrator, guided by real-time capacity profiling, re-splits and reconfigures workloads on demand to meet QoS and privacy constraints.
@@ III-A Reference Architecture
Our approach dynamically adapts split inference to fluctuating conditions while maintaining strict QoS and privacy requirements by combining these modules. The next subsections formalize the system model, define constraints, and describe the orchestration workflow for LFMs in detail.
III-B Notation and System Model
We define key terminologies and orchestration concepts that underlie our adaptive split-inference framework as follows.
@@ III-C Orchestration Workflow
Input: (i) Initial partitioning $\{S_{1},\dots,S_{P}\}$, (ii) baseline mapping $d_{0}$, (iii) monitoring intervals $\Delta t$.
       Collect environment metrics $\mathbf{E}(t)$ via Monitoring & CP.
       if (trigger condition is met, e.g., high latency, node overload, etc.) then
             Evaluate feasible mappings $\{d^{\prime}\}$ given current partitions.
             Determine best mapping $\hat{d}=\underset{d^{\prime}}{\operatorname{argmin}}\,\mathcal{C}(d^{\prime})$.
             if $\hat{d}\neq d_{t}$ then
                   Broadcast reconfiguration to all affected nodes via RB.
                   Update partitioning and mapping: $d_{t+\Delta t}\leftarrow\hat{d}$.
             end if
       Resume inference under current assignment $d_{t+\Delta t}$.
end for
Algorithm 1 Adaptive Split Orchestration Workflow
@@
III-D Privacy and Security Considerations

### v2

@@ Abstract
Large Foundation Models (LFMs), including multi-modal and generative models, promise to unlock new capabilities for next-generation Edge AI applications. However, performing inference with LFMs in resource-constrained and heterogeneous edge environments, such as Multi-access Edge Computing (MEC), presents significant challenges for workload orchestration due to time-varying network, compute, and storage conditions. In particular, current split inference strategies, which partition LFM layers across nodes, are not designed to adapt to fluctuating workloads, dynamic bandwidth conditions, or evolving privacy constraints in high-utilization MEC environments. In this work, we propose a novel adaptive split inference orchestration framework that elevates both the placement and partitioning of LFM layers to runtime-tunable variables. Specifically, our framework enables real-time, quality-of-service (QoS)-aware management of inference workloads by extending conventional orchestrators with three key services: (1) Capacity-aware workload distribution, which continuously profiles node resources and selects an optimal subset of MEC nodes; (2) Dynamic partition migration, which transparently relocates pre-cut LFM segments in response to changes in utilization or network conditions; (3) Real-time reconfiguration, which dynamically re-splits LFM layers to balance latency, throughput, and privacy. We formalize the joint placement-partitioning problem, outline a reference architecture and algorithmic workflow, and discuss applicability in representative smart city, V2X, and industrial edge scenarios.
1 Introduction
Inference, the real-time forward execution of a trained model, is typically less compute-intensive than model training and may require substantial resources for layer activation, memory, and model storage [1]. This holds particularly for Large Foundation Models (LFMs) such as transformer-based large language models (LLMs) [2], e.g., Llama [3] and Qwen [4] models. Performing inference efficiently on the edge remains challenging, especially in Multi-Access Edge Computing (MEC) environments with limited and heterogeneous compute resources [5].
Distributed Split Inference (DSI) [6] has emerged as a promising approach to mitigate this problem. This strategy aims to partition an LFM into multiple segments that are executed sequentially across different nodes.
@@ 1 Introduction
DSI strategies have typically employed static splits of inference workloads, where some computation is executed locally, while heavier computation tasks are outsourced, alleviating the computational burden on the client device. Such splits are mostly predetermined before execution and therefore lack adaptability to dynamic and heterogeneous operational conditions, such as fluctuating network reliability, changing node utilization, or intermittent connectivity. Consequently, these approaches produce suboptimal performance, compromising latency, resource utilization, and quality of service (QoS) guarantees, especially in mission-critical or latency-sensitive applications such as those found in financial services, industrial manufacturing, retail operations, and logistics [7]. This problem becomes even more acute in resource-constrained and heterogeneous edge environments, where multiple users rely on accessing shared edge compute resources, and where data privacy regulations (e.g., GDPR [8]) often restrict offloading computations to the cloud. For instance, executing a 7B parameter LLM may incur 25 ms on an NVIDIA RTX A6000 edge node but over 250 ms on an NVIDIA Jetson accelerator. Meanwhile, backhaul latency can oscillate between sub-1 ms (mmWave) and 30 ms (congested Wi-Fi) [9, 10]. Such volatility renders any a priori static split untenable.
Contemporary orchestration frameworks such as Kubernetes, Ray Serve, InferLine, and KubeEdge excel at container or micro-batch scheduling but treat AI models as black boxes, lacking mechanisms for runtime layer re-partitioning or privacy-aware placement [11].
Further, recent works, e.g., on jamming resilience for Split Federated Learning with LLMs[12], Federated Split Learning for satellite–terrestrial networks [13], and on Split Federated Mutual Learning (SFML) for traffic classification [14], address training-time collaboration, yet still assume fixed split points. Consequently, the key problem of joint, model-aware partition and placement under dynamic edge conditions remains open.
@@
In this work, we address this problem by introducing an adaptive split inference orchestration framework, extending existing workload orchestration systems with domain-specific capabilities that are specifically tailored for LFMs, such as (multi-modal) LLMs. We introduce the following capabilities by leveraging the modular architecture of these models:
@@
    Redistribution of split LFM partitions across connected edge nodes to dynamically optimize resources under changing conditions.
@@
Our framework operates on the computational graph of the LFM itself, allowing decisions at the granularity of individual transformer blocks. This enables QoS-driven re-splitting that commodity orchestrators cannot express. Furthermore, since our solution emphasizes split inference, privacy can be implemented as an additional feature at no cost if sensitive LFM layers can be executed locally, which makes reverse engineering data from model weights significantly more challenging for attackers [15]. Through this approach, we establish a foundation for privacy-preserving, real-time, and QoS-aware AI inference in edge networks, aligning with key 6G objectives of seamless connectivity, low inference latency, and intelligent edge resource management [16]. Importantly, our framework elevates privacy guarantees by localizing sensitive computations, thereby reducing regulatory exposure and aligning with emerging compliance standards.
Our contributions to the Edge AI landscape are multifold:
  • a reference architecture for adaptive orchestration of distributed inference workloads;
  • a method for dynamic redistribution of model segments to accommodate fluctuating compute resources and connectivity, and;
  • establish mechanisms for real-time, Service-Level Agreement (SLA)-compliant partitioning that align inference execution with QoS targets.
2 Background and Related Work
2.1 Large Foundation Models at the Edge
@@
Next-generation 6G-enabled services, for example in dense urban environments, will need to support a multitude of AI-driven applications underpinned by LFMs [17, 18]. However, deploying such large models typically requires significant computational resources and raises privacy concerns when handling sensitive data, making their adoption particularly challenging for inference in edge environments [19, 20]. Further, as organizations strive to keep data on-premise (e.g., for regulatory compliance such as GDPR [8]), optimizing distributed inference becomes a critical enabler of low-latency and privacy-preserving AI services [21, 1].
@@
For that, next-generation networks make use of MEC infrastructures [22], which embed computing resources directly within the network. While this brings compute closer to end-users, a single MEC-enabled base station can quickly become saturated by various inference workloads such as from smart city and crowd management, personalized user applications, or industrial applications [23]. Looking ahead, 6G is envisaged to evolve beyond a mere network infrastructure upgrade into a trustworthy [24] and intelligent workload orchestration system, enabling distributed, LFM-based AI services that seamlessly shift computation across user devices, edge, and cloud as network and compute conditions change [25, 26].
2.2 Challenges in AI Workload Orchestration
@@
The current industry norm has been to integrate general-purpose orchestration platforms (e.g., Kubernetes or proprietary MEC orchestrators), which facilitate application deployment and scaling but were not designed for challenges in inference scaling of modern LFM architectures, including fine-grained model partitioning and real-time, potentially hardware-accelerated inference optimization [27]. Existing solutions thus primarily target stateless services or relatively simple microservices, neglecting the unique requirements of edge-based AI pipelines, particularly for dynamic model splitting, QoS-driven scheduling, and adaptive resource reallocation [28].
@@
In particular, in Edge AI scenarios involving LFMs, traditional workload orchestration platforms fall short of meeting Edge AI-centric inference demands as they lack mechanisms to dynamically redistribute or reconfigure large model partitions based on real-time changes in network conditions, node utilization, or connectivity [29]. While model-serving stacks such as Triton, InferLine, Ray Serve, and MLC-Serve introduce batching or replica autoscaling, they still treat neural networks as opaque binaries and cannot re-partition a model graph at runtime. Edge extensions like KubeEdge and OpenYurt inherit the same pod-level abstraction, leaving the joint problem of runtime layer splitting and placement, especially under privacy and QoS constraints, unaddressed (see Table 1).
Table 1: Comparison of capability coverage across mainstream workload orchestrators / serving stacks and our proposed Adaptive Split Orchestrator. ✓ indicates native support, ✗ indicates the capability is absent, and entries in parentheses denote the mechanism (e.g., pods, replicas). Our framework simultaneously (i) reasons over the layer graph, (ii) adapts split points and placements at runtime, (iii) schedules under explicit QoS/SLA targets, and (iv) enforces privacy-aware layer placement.
Orchestrator Layer-Aware Graph Runtime Split Placement Adapt. QoS- Aware Privacy Support
Kubernetes (pods)
KubeEdge (edge pods)
Ray Serve (replicas) (latency tiers)
NVIDIA Triton (inst. groups) (batch/queue)
InferLine (replicas) (tail-latency)
EdgeShard [30] (manual) (manual) N/A
Ours (runtime) (runtime) (SLA-aware) (layer scope)
As a result:
@@
  • Resource utilization becomes imbalanced, either overloading certain nodes or leaving others underutilized, leading to missed SLAs.
@@
2.3 Distributed and Adaptive Split Inference
To alleviate computational demands, distributed split inference (DSI) [34] has emerged as an approach within Edge AI, which partitions a model across different compute locations (e.g., client device, MEC node, cloud) to balance local processing with remote offloading. Here, early, lightweight, or privacy-sensitive model layers are often chosen to be executed locally on-device or on trusted MEC nodes, extracting compact feature maps to reduce data transmission overhead. Subsequent layers are then offloaded to remote cloud environments and processed by more powerful servers [34, 35].
@@ 2.3 Distributed and Adaptive Split Inference
In addition to inference, recent split learning approaches illustrate the state of today’s limitations from the model training point of view. R-SFLLM [12] freezes an LLM split immediately after the embedding layer of a transformer so that raw tokens never leave the device, and then analyses jamming resilience for that single layout. FedSL-LSTM [13] for satellite–terrestrial anomaly detection pre-divides each LSTM into client- and server-side subnetworks before training begins. The interface remains unchanged to keep gateway compute low and ground timing predictable. SFML [14] adopts a fixed three-segment CNN in which only lightweight “head+tail” layers stay on edge routers while the compute-heavy core runs in the cloud, simplifying privacy guarantees but eliminating runtime flexibility. Across all three studies, split points are determined a priori and never migrated, leaving questions of dynamic re-splitting, QoS-driven placement, and privacy-aware adaptation unanswered, precisely the gaps our adaptive orchestrator addresses. Table 2 summarizes these representative frameworks, highlighting their fixed partitioning schemes and the lack of adaptive capabilities.
Table 2: Comparison of representative split learning frameworks with respect to their application context, model partitioning strategies, and (lack of) adaptivity. While each approach adopts fixed split points to simplify deployment or privacy guarantees, none supports dynamic re-partitioning or runtime migration, highlighting the need for adaptive orchestration in real-world, QoS-sensitive Edge AI scenarios.
R-SFLLM FedSL-LSTM SFML
Application Jamming-resilient FL with LLMs over 6G wireless links LSTM anomaly detection in satellite–terrestrial integrated networks Encrypted-traffic classification on edge routers + cloud
Split Layer After the embedding block; attention layers off-device First LSTM layers on gateways; remainder on ground server Lightweight head + tail on router; CNN core in cloud
Adaptivity Static (single two-cut split, never moved during training) Static (single two-cut split, never moved during training) Static (single three-cut split, never moved during training)
Thus, in standard implementations, orchestrators cannot dynamically decide to offload additional LFM layers to another edge node or re-split the network, leaving significant performance and reliability gains unrealized. Note that this problem does not originate from within the network itself as current 5G and future 6G architectures natively implement adaptive resource management strategies for various services and network slices [37]. In current deployments, effectiveness thus remains limited by the absence of real-time dynamic orchestration policies tailored explicitly to modern AI workloads and complex LFM deployment scenarios. This is especially problematic for heterogeneous compute nodes, which complicates uniform deployment strategies. Thus, a one-size-fits-all static partitioning rarely works, as local workloads, performance constraints, and available resources differ from one site to another [19].
2.4 Key Design Goals for Adaptive LFM Split Inference
@@ 3 Proposal
3.1 Reference Architecture
Figure 1: Reference architecture of the proposed adaptive split inference orchestration. Sub-split models (S1, S2, S3) are deployed across edge/cloud nodes, while a central orchestrator, guided by real-time capacity profiling, re-splits and reconfigures workloads on demand to meet QoS and privacy constraints. A corresponding workflow diagram of our proposed Algorithm 1 is given in Figure 2.
@@ 3.1 Reference Architecture
Our approach dynamically adapts split inference to fluctuating conditions while maintaining strict QoS and privacy requirements by combining these modules. Although Kubernetes already offers node-level telemetry, a pod scheduler, and ReplicaSet roll-outs, it treats the AI model as an opaque container. Instead, our Monitoring & CP module augments the standard metrics stream with layer-granular latency, activation size, and privacy tags. The Adaptive Orchestrator then optimizes both placement and split boundaries, whereas Kubernetes can only relocate whole pods. Finally, the Reconfiguration Broadcast disseminates on-the-fly–generated weight shards and graph rewiring commands—operations that cannot be expressed through a deployment update or ConfigMap patch. Collectively, these extensions lift orchestration from container granularity to LFM-graph granularity, which is essential for runtime re-splitting under QoS and privacy constraints.
The next subsections formalize the system model, define constraints, and describe the orchestration workflow for LFMs in detail.
3.2 Notation and System Model
We define key terminologies and orchestration concepts that underlie our adaptive split inference framework as follows.
@@ 3.3 Orchestration Workflow
    Input: (i) Initial partitioning $\{S_{1},\dots,S_{P}\}$, (ii) baseline mapping $d_{0}$, (iii) monitoring intervals $\Delta t$, (iv) trigger-threshold vector $\Theta=\{L_{\max},U_{\max},B_{\min},T_{\text{cool}}\}$
    Set $t_{\mathrm{last}}\leftarrow-\infty$.
           Collect environment metrics $\mathbf{E}(t)$ via Monitoring & CP.
[…]
                 Evaluate feasible mappings $\{d^{\prime}\}$ given current partitions.
                 Determine best mapping $\hat{d}=\underset{d^{\prime}}{\arg\min}\,\mathcal{C}(d^{\prime})$.
                 if $\hat{d}\neq d_{t}$ and $t-t_{\mathrm{last}}\geq T_{\text{cool}}$ then
                       Broadcast reconfiguration to all affected nodes via RB.
                       $t_{\mathrm{last}}\leftarrow t$;  $d_{t+\Delta t}\leftarrow\hat{d}$.
                 end if
           Resume inference under current assignment $d_{t+\Delta t}$.
    end for
    Algorithm 1 Adaptive Split Orchestration Workflow
@@
Additional Trigger Conditions and Decision Logic. Table 3 summarizes the runtime metrics that feed the function ShouldReconfigure(E(t),$\Theta$) in Algorithm 1. The orchestrator invokes a reconfiguration if any of the following holds for a monitoring window of length $\Delta t$:
  1. 1.
    Latency threshold. The exponentially weighted moving average (EWMA) of end-to-end inference latency $\bar{L}(t,\Delta t)$ exceeds $L_{\max}$, i.e. $\bar{L}(t,\Delta t)>L_{\max}$.
  2. 2.
    Utilisation threshold. The maximum GPU or CPU utilization across all nodes exceeds $U_{\max}$, i.e., $\max_{n\in\mathcal{N}}U_{n}(t)>U_{\max}$.
  3. 3.
    Bandwidth drop. The minimum available bandwidth across any active edge link drops below $B_{\min}$, i.e., $\min_{(i,j)\in\mathcal{L}}B_{ij}(t)<B_{\min}$.
  4. 4.
    Privacy policy violation. A new inference request is tagged with the identifier privacy=high, but the current partitioning would route raw features through an untrusted node.
Table 3: Monitored metrics and default trigger thresholds.
Metric Symbol Default value
EWMA latency $L_{\max}$ 150 ms
GPU/CPU utilisation $U_{\max}$ 0.85
Available link bandwidth (edge $\to$ edge) $B_{\min}$ 50 Mbps
Time-to-reconfigure cool-down $T_{\text{cool}}$ 30 s
If multiple triggers fire simultaneously, the system first attempts placement migration. If that cannot meet all constraints, the split-revision module is invoked. Reconfigurations are rate-limited by $T_{\text{cool}}$ to prevent thrashing. The full control loop of our adaptive orchestration algorithm is illustrated in Figure 2, detailing its monitoring, decision, and reconfiguration stages.
Initialize: deploy baseline split $d_{0}$, $t_{\rm last}\leftarrow-\infty$ Collect metrics $\mathbf{E}(t)$ Should Reconf? $(\mathbf{E},\Theta)$ Resume under $d_{t+\Delta t}$ Eval feasible $d^{\prime}$ maybe ReSplit $\hat{d}\leftarrow\arg\min_{d^{\prime}}\mathcal{C}(d^{\prime})$ Reconf & Cool-down? Resume under $d_{t+\Delta t}$ Broadcast $d_{t+\Delta t}\leftarrow\hat{d}$ $t_{\rm last}\leftarrow t$ Resume under $d_{t+\Delta t}$ Wait $\Delta t$ $t\!\leftarrow\!t+\Delta t$ NoYesNoYes
Figure 2: Control-flow diagram of the adaptive split orchestration loop described in Algorithm 1. The orchestrator periodically monitors environment metrics and triggers reconfiguration decisions when QoS thresholds or privacy constraints are violated. Feasible placements are evaluated, and, if no cool-down limit is active, a new mapping is broadcast to all nodes.
@@
3.4 Privacy and Security Considerations